Solve each problem and provide the complete solution.
note: topic is all about Gravitational force

What is the gravitational potential energy of a girl who has a mass of 50 kg and is standing on the edge of a diving board that is 5m above the water?
1
Expert's answer
2020-11-30T14:57:12-0500

By the definition of the gravitational potential energy, we have:


"E_p=mgh=50\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot 5\ m=2450\ J."

Answer:

"E_p=2450\ J."